1. A capacitor consists of two 4.8-cm-diameter circular plates separated by 1.0 mm. The plates are charged to 170 V , then the battery is removed.

How much work must be done to pull the plates apart to where the distance between them is 2.0 mm? (Express your answer in joules.)

2. A 2.5 mm -diameter sphere is charged to -4.4 nC . An electron fired directly at the sphere from far away comes to within 0.30 mm of the surface of the target before being reflected.

At what distance from the surface of the sphere is the electron's speed half of its initial value? ANSWER IN METERS.
Answer #1

Capacitance C = 8.854×10^(-12) × 3.14× (0.024)^2/0.001

= 1.601 ×10^(-11)

Now energy stored E = 1/2× C × V^2

= 0.5×1.601×10^(-11) ×170×170 = 2.314×10^(-7) J

Whenc plate displace i.e distance become double, capacitance become half and hence energy stored will become half. Hence work done will be half of the energy stored i.e w = 1.157 ×10^(-7) J

Qus 2----